### v2

# Exploring Kolmogorov-Arnold Networks for Interpretable Time Series Classification

Abstract
Time series classification is a relevant step supporting decision-making processes in various domains, and deep neural models have shown promising performance. Despite significant advancements in deep learning, the theoretical understanding of how and why complex architectures function remains limited, prompting the need for more interpretable models. Recently, the Kolmogorov-Arnold Networks (KANs) have been proposed as a more interpretable alternative. While KAN-related research is significantly rising, to date, the study of KAN architectures for time series classification has been limited. In this paper, we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification on the UCR benchmark. More specifically, we look at a) how reference architectures for forecasting transfer to classification, at the b) hyperparameter and implementation influence on the classification performance in view of finding the one that generalizes best on the classification benchmark, the c) complexity trade-offs and d) interpretability advantages. Our results show that (1) Efficient KAN outperforms MLP in performance and computational efficiency, showcasing its suitability for tasks classification tasks. (2) Efficient KAN is more stable than KAN across grid sizes, depths, and layer configurations, particularly with lower learning rates. (3) KAN maintains competitive accuracy compared to state-of-the-art models like HIVE-COTE2, with smaller architectures and faster training times, supporting its balance of performance and transparency. (4) The interpretability of the KAN model aligns with findings from SHAP analysis, reinforcing its capacity for transparent decision-making.
Index Terms:
classification, time series, Kolmogorov Arnold networks, multilayer perceptrons
I Introduction
Time series classification is a relevant step supporting decision-making processes in various domains. In healthcare for example, time series classification enables the diagnosis and monitoring of conditions by analyzing patterns in physiological signals such as electrocardiograms or brain activity [1]. In finance, it enhances risk management and fraud detection by classifying trading behaviors and transaction patterns [2]. By classifying activities based on sensor data in human activity recognition, it supports personalized recommendations and safety monitoring [3]. In remote sensing, it contributes to environmental monitoring and land use classification by categorizing satellite or aerial imagery data over time [4]. In these fields, accurate time series classification plays a central role in advancing predictive and analytic capabilities that drive critical outcomes.
Unlike time series forecasting, which aims to predict future data points, time series classification aims to provide accurate interpretation and categorization of temporal data sequences by assigning a categorical label, class label, to each sequence [5]. Following a well-accepted survey [6] that revealed the lack of investigation of deep learning techniques for time series classification, significant progress has been achieved in related research with hundreds of new works, some of them surpassing the non-deep state of the art [7]. According to the most recent and extensive benchmarking work on 100+ datasets from the UCR archive, the most extensive benchmark for time series classification, the state of the art F1 score of 0.886 is achieved by Hive-Cote2 [8] [9], a large ensemble with considerable training time. The non-deep and non-hybrid classification methods that are more explainable at both feature engineering and decision making steps achieve an F1 of up to 0.869 while the deep architectures reach around 0.88 F1 score. Foundation models for time series, based on transformers, are yet to be studied on cross-domain benchmarks such as UCR [10].
Despite the disruptive advancements introduced by breakthroughs in machine vision and natural language processing over the last decades, the theoretical understanding of why and how complex deep architectures function has lagged behind [11]. This has prompted researchers to step back and reconsider fundamental and mathematically simpler architectures, such as MLPs [12]. Furthermore, the influence of the academic scientific community in this new era of AI is decreasing [13], while non-AI scientific communities are concerned with the interpretability of deep models [11]. Recently, the Kolmogorov-Arnold Networks (KANs) [14] have been proposed as addressing the limitations of traditional neural networks. KANs have demonstrated to be interpretable, even enabling symbolic regression, and have comparable performance with MLPs on small scale and science tasks with shallower architectures. Additional advantages have been explored for data fitting and solving partial differential equations [15].
The main critique of the original KAN work is concerned with the fairness of the comparison, triggering substantial research into better understanding their overall performance beyond the initial small scale scientific tasks. Subsequently it has been shown that their performance advantage in non-scientific ML tasks such as vision, natural language and audio processing does not hold [16]. Additional work reveals their reduced effectiveness on functions with noise [17] while other works investigate the suitability of replacing the spline functions with wavelets [18]. The performance of KANs vs MLPs on graph learning tasks is compared in [19]. Their preliminary results reveal that while KANs are on-par with MLPs in classification tasks, they seem to have a clear advantage in the graph regression tasks.
While KAN-related is significantly rising, new scientific works published almost weekly, to date, the study of KAN architectures for time series has been limited. Very early forecasting studies encompass Temporal KANs (TKANs) [20] and Temporal Kolmogorov Arnold Transformer [21] for notional trades, mixture of experts [22] for various domains including traffic and weather, satellite traffic [23] and integrating convolutional layers with KANs to improve time-series forecasting [24]. Furthermore, with respect to time series classification, studies on their robustness are emerging [25].
In this paper we aim to conduct a comprehensive and robust exploration of the KAN architecture for time series classification on the UCR benchmark. More specifically, we look at a) how the existing architectures for forecasting [23] transfer to classification, at the b) hyperparameter and implementation influence on the classification performance in view of finding the one that performs best on the selected benchmark, the c) complexity trade-ffs and d) interpretability advantages. The contributions of this paper are as follows.
  • A study on the suitability of KAN architectures for classification tasks on the UCR dataset. The study first investigates the feasibility of transferring existing architectures designed for forecasting followed by finding the most suitable architecture for classification.
  • A hyperparameter impact analysis of two KAN implementations, analyzing how variations in grid size, network depth, and node configurations impact classification performance. The analysis leads to finding the best performing KAN configuration for the considered problem.
  • The performance and computational complexity comparison of the original KAN, the Efficient KAN implementation, and Multi-Layer Perceptrons (MLPs) on time series classification tasks reveals EfficientKAN’s superior stability across grid sizes depths and layer configurations.
  • We confirm KAN’s interpretability by diving deeper into the learnt feature importance and activations functions vs SHAP importance.
The paper is structured as follows. We discuss related work in Section II. Section III outlines the problem statement, followed by Section IV detailing methodological aspects. A comprehensive analysis of the results is presented in Section V. Lastly, the paper concludes with Section VI.
II Related work
Kolmogorov-Arnold Networks (KANs) emerged as an innovative alternative to traditional Multi-Layer Perceptrons (MLPs), inspired by the Kolmogorov-Arnold representation theorem. KANs were shown to outperform MLPs in both accuracy and interpretability on small-scale scientific tasks. Their design, which models complex functions and patterns with fewer parameters, demonstrated potential to aid in mathematical and physical discoveries [14]. Building on these findings, subsequent research [15] further developed KANs to bridge the gap between artificial intelligence and scientific research. By allowing KANs to identify relevant features, reveal modular structures, and discover symbolic formulas, KAN 2.0 [15] introduced a bidirectional approach that not only incorporates scientific knowledge into KANs but also enables KANs to extract interpretable scientific insights from data. However, KANs have been shown to be highly sensitive to noise, prompting the introduction of oversampling and denoising techniques, such as kernel filtering with diffusion maps, to mitigate noise effects[17]. To improve the handling of noisy data, architectural extensions have additionally explored the replacement of splines with wavelets [18], enhancing KAN’s robustness and adaptability in broader applications.
KANs and MLPs were also investigated beyond scientific datasets in various domains in a controlled study with consistent parameters and FLOPs [16]. While MLPs outperformed KANs in most areas, KANs retained a distinct advantage in symbolic formula representation due to B-spline activation functions. Replacing MLPs’ activations with B-splines improved their performance, suggesting KAN-inspired enhancements for MLPs. However, KANs faced memory stability issues in continual learning, requiring specialized tuning or hybrid models. In graph learning tasks [19], experiments on node classification, graph classification, and graph regression datasets indicated that KANs are on par with MLPs in classification but exhibit an advantage in graph regression tasks.
In a real-world satellite traffic forecasting task, KANs achieved comparable or superior accuracy to MLPs while using fewer parameters, showcasing their potential in predictive analytics [23]. This application framed forecasting as a supervised learning problem with specific input-output mappings across time steps, using a GEO satellite traffic dataset. The success of KANs in this setting prompted specialized variants, such as Temporal KAN (T-KAN) and Multivariate Temporal KAN (MT-KAN) [26]. T-KAN targets univariate time series, capturing nonlinear relationships with symbolic regression, while MT-KAN model dependencies between multiple variables for improved accuracy in multivariate settings. Further advancements, like Temporal Kolmogorov-Arnold Networks (TKANs), incorporated LSTM-inspired memory layers, excelling in tasks such as cryptocurrency trading volume forecasting. Temporal Kolmogorov-Arnold Transformer (TKAT) added self-attention mechanisms, outperforming conventional transformers in interpretability and precision in multivariate time series forcasting [20, 21]. Signature-Weighted KANs (SigKAN) extended these innovations by integrating path signatures, making them robust for market volume prediction [27]. The Reversible Mixture of KAN Experts (RMoK) model, another KAN-based approach, introduced a mixture-of-experts structure to assign variables to KAN experts, achieving strong performance in time series forecasting tasks by leveraging temporal feature weights to explain data periodicity [22].
To enhance time series classification robustness, hybrid models that combine KAN and MLP architectures were further explored [25]. The study employed Efficient KAN as the primary implementation of KAN rather than the original implementation. Two hybrid configurations, KAN MLP (KAN with an MLP as the final layer) and MLP KAN (MLP with a KAN as the final layer), were tested across the UCR datasets, and both hybrids achieved performance comparable to traditional KAN and MLP models. Notably, MLP KAN demonstrated increased resilience against adversarial attacks, with this robustness attributed to the lower Lipschitz constant of KAN layers, suggesting that combining KAN with MLP structures can strengthen adversarial resistance in neural networks, thus opening possibilities for more secure and reliable models. Rather than focusing on robustness and hybrid architecture, this study aims at understanding in depth the performance and interpretability of KANs for time series classification.
III Problem Statement
In this paper we analyze the potential of KAN for solving univariate time-series classification problems. A univariate time series $X$ consists of a sequence of observations $[x_{1},x_{2},\ldots,x_{T}]$, where each $x_{t},t\in\{1,2,...,T\}$ represents a value observed at time $t$ representing an ordered set of real values [6].
A class label $Y$ is a categorical variable associated with a time series $X$, indicating the class that the time series belongs to. The number of distinct classes is denoted by the cardinality $C$ of $Y$, meaning $Y$ can take a value from the set $\{y_{1},y_{2},\ldots,y_{C}\}$. When $C=2$, this corresponds to a binary classification problem, while $C>2$ represents a multiclass classification problem. As shown in Fig. 1, where $C=3$, each plot corresponds to a different class, illustrating the distinct patterns for $y_{1}$, $y_{2}$, and $y_{3}$.
A dataset $D=\{(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{N},Y_{N})\}$ consists of pairs $(X_{i},Y_{i})$, where $X_{i},i\in\{1,2,3...,N\}$ represents a univariate time series, and $Y_{i}$ is the corresponding class label.
The task of time series classification involves training a classifier on a dataset $D$ to map each input $X_{i}$ to a probability distribution across the possible class labels [6].
$f(D)\approx Y$ (1)
Specifically, instead of assigning a single class label to each time series, the classifier outputs a probability distribution across all possible classes. That is, for each time series $X_{i}$, the classifier predicts a vector of probabilities $P(Y_{i}=y_{c}\mid X_{i})$ for each class $c\in\{1,2,\dots,C\}$.
The classification function $f$ can be realized through a multitude of exiting approaches [7], however we investigate the recently proposed Kolomogrow Arnold Network (KAN) [14] promising increased interpretability and potential computational benefits compared to standard Multilayer Perceptron (MLP).
Figure 1: Example of time series data plotted by class, with time steps on the x-axis and observed values $x_{t}$ on the y-axis. Figure depicts SmoothSubspace dataset from the UCR repository.
III-A KAN for time series classification
The architecture of a KAN is depicted in Fig. 2. It can be seen from the figure that, for an input instance $X_{i}$, the network outputs $Y_{i}$ label assignment probability vector. $Y$ is learnt from $D$ as a composition of $L$ appropriate univariate functions $\Phi$, as depicted in Fig. 2 and formalized in Eq. 2, where $L$ stands for the number of layers in the network.
[…]
Figure 2: KAN architecture.
The structure of a KAN can be represented as $[n_{1},…,n_{L+1}]$, where $L$ signifies the total number of layers in the KAN. A deeper KAN can be thus formulated through the composition of $L$ layers as:
$Y=\text{KAN}(X)=(\Phi_{L}\circ\Phi_{L-1}\circ\dots\circ\Phi_{1})X$ (2)
Each layer of a KAN is represented by a matrix where each entry is an activation function. If there is a layer with $d_{in}$ nodes and its neighboring layer with $d_{out}$ nodes, the layer can be represented as a $d_{in}\times d_{out}$ matrix of activation functions (Eq. 4).
$\Phi=\{\phi_{q,p}\},\;p=1,2,\dots,d_{in},\;q=1,2,\dots,d_{out}$ (3)
Unlike traditional MLPs, where activation functions are applied at the nodes themselves, KAN places them at the edges between the nodes. KAN employs the SiLU activation function in combination with B-splines to enhance its expressiveness. This setup allows the edges to control the transformations between layers, while the nodes perform simple summation operations.
$\phi(x)=w_{b}silu(x)+w_{s}\,\text{spline}(x)$ (4)
where SiLU is defined as $\text{silu}(x)=x/(1+e^{-x})$. This activation function allows for smooth, non-linear transformations, which helps the network capture complex patterns in time series data more effectively.
B-splines are smooth polynomial functions of degree (order) $k$, which approximate data using control points. The order $k$ controls the smoothness of the spline. Commonly, $k=3$ is used for cubic splines. Each spline operates over a defined grid $G$, which divides the input space into smaller intervals. The grid is determined by a set of grid points that define the segments of the spline. Increasing the number of grid points, increases spline resolution, enabling more fine grained approximation of the underlying univariate function.
A B-spline of order $k$ requires $G+k$ basis functions to represent the spline over the grid. For each input (node in a layer), the evaluation of a B-spline of order $k$ thus involves computing $G+k-1$ basis functions and performing a weighted sum with the corresponding control points.
$\text{spline}(x)=\sum_{i=0}^{G+k-1}c_{i}B_{i}(x)$ (5)
To better understand the performance of the models evaluated, it is crucial to explain their decision-making processes. Although black-box models achieve high accuracy, they often lack interpretability. KANs offer a shift in the interpretability paradigm by allowing visualization of the model structure along with its learned B-splines as shown in Fig. 2. Interpretability is further enhanced by model pruning that simplifies visual representation by removing less important connections within the model, which were determined through their significance scores. These scores identify critical connections, which highlight the most important features for the decision of the model. This approach provides insights into how KANs process data and make predictions by focusing on the most influential connections within the structure of the model.
III-A1 KAN learnable parameters
The number of learnable parameters in a KAN layer is determined by its architecture, which includes contributions from the B-spline control points, shortcut path weights, B-spline weights, and bias terms. As discussed in [16], the number of learnable parameters for each KAN layer is:
$\text{Parameters}=(d_{in}\cdot d_{out})\cdot(G+k+3)+d_{out}$ (6)
Where $d_{in}$ and $d_{out}$ represent the input and output dimensions of the layer. This is consistent with the findings in the [14], which also emphasize the role of learnable control points and weights associated with the spline functions.
III-A2 KAN FLOPs computation
The number of FLOPs in a Kolmogorov–Arnold Network (KAN) is highly dependent on the specific implementation and the way the network is compiled on various processor architectures. Different hardware platforms optimize these computations differently, leading to variations in practical FLOPs performance. However, we adopt the theoretical FLOPs computation model as outlined in [16] for consistency and comparison purposes.
The total FLOPs for a KAN layer come from three parts: the B-spline transformation, the shortcut path, and the merging of the two branches. Using the De Boor-Cox iterative formulation for B-splines, the FLOPs for one KAN layer in the official KAN implementation is given by [16]:
FLOPs=FLOPs of non-linear $\begin{split}\text{FLOPs}=&\ \text{FLOPs of non-linear function}\cdot d_{\text{in}}\\ &+(d_{\text{in}}\cdot d_{\text{out}})\cdot\left[9\cdot k\cdot(G+1.5\cdot k)\right.\\ &\left.+2\cdot G-2.5\cdot k+3\right]\end{split}$ (7)
FLOPs of one forward pass $FP$ through a network with uniform hidden layers of size $d_{in}\cdot d_{out}=M\cdot M$ are calculated as in Eq. 8:
$\begin{split}N_{FP}=&\ \text{FLOPs of non-linear function}\cdot T\\ &+T\cdot M\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\ +2\cdot G-2.5\cdot k+3\big{]}\\ &+(L-2)\cdot\big{(}\text{FLOPs of non-linear function}\cdot M\\ &\quad\ +M^{2}\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\quad\ +2\cdot G-2.5\cdot k+3\big{]}\big{)}\\ &+\text{FLOPs of non-linear function}\cdot M\\ &+M\cdot C\cdot\big{[}9\cdot k\cdot(G+1.5\cdot k)\\ &\quad\ +2\cdot G-2.5\cdot k+3\big{]}\end{split}$ (8)
III-B MLP time series classification
A Multilayer Perceptron (MLP) is a type of feed-forward artificial neural network, used in various machine learning tasks, including time series classification. MLPs are inspired by the universal approximation theorem, which asserts that such networks can approximate any continuous function given sufficient depth (number of layers) and width (number of neurons per layer).
[…]
Figure 3: A fully connected MLP architecture.
The architecture of a MLP is depicted in Fig. 3. The MLP consists of multiple layers of nodes (neurons): input layer with $T$ neurons, $K$ hidden layers (layers which are not directly connected to the environment are called hidden [28]) and output layer, consisting of $C$ output neurons.
Each neuron in a layer processes the input it receives from the previous layer by calculating a weighted sum of the inputs, adding a bias term, and then passing the result through an activation function. For a neuron with $n$ inputs, its output is given by the Eq.:
$neuron=a\left(\sum_{i=1}^{n}weight_{i}\cdot input_{i}+bias\right)$ (9)
where $a$ represents the activation function. A commonly used activation function in MLPs is the Rectified Linear Unit (ReLU), defined as $\text{ReLU}(z)=\max(0,z)$.
This structure allows MLPs to model complex, non-linear relationships between inputs and outputs. In time series classification, the input to the MLP $X$, is a fixed-length vector representing the time series data. The network learns to map this input vector to a probability distribution across the possible class labels, where each neuron $p_{c},c\in\{1,2,\dots,C\}$ in the final output layer represents the probability of the input belonging to a specific class $y_{c}$.
III-B1 MLP learnable parameters
The number of learnable parameters in an MLP is determined by the connections between neurons across layers. For a fully connected layer with $d_{in}$ input neurons and $d_{out}$ output neurons, the number of learnable parameters [16] is:
$Parameters=(d_{in}\texttimes d_{out})+d_{out}$ (10)
III-B2 MLP FLOPs computation
In a fully connected MLP, each connection between two neurons $wx+b$ performs a weighted sum $w\cdot x$ and adds bias $b$, resulting in 1 multiplication and 1 addition. Thus, each connection requires 2 FLOPs.
For a neuron with $d_{\text{in}}$ inputs, calculating the output requires $2\times d_{\text{in}}$ FLOPs (for weights) and 1 FLOP for the bias, resulting in a total of $12\times d_{\text{in}}+$ 1 FLOPs. If there are $d_{\text{out}}$ output neurons, the total FLOPs for the layer is $d_{\text{out}}\times(2\times d_{\text{in}}+1)$.
We consider fully connected MLP which has input layer with $T$ neurons, $K$ hidden layers, with $M$ neurons in each, and $C$ neurons in output layer. FLOPs for forward propagation are therefore calculated as
$N_{FP}=(M+2MT)+(K-1)(M+2M^{2})+(M+2MC)$ (11)
IV Methodology
In this section we elaborate on the methodology adopted for this study ($https://github.com/irina-b1/KAN_{T}S$). First we provide considerations on data and preprocessing, followed by reference model configurations, training process, hyperparamater, complexity and interpretability analysis.
IV-A Data and preprocessing
In this study, the UCR (University of California, Riverside)[29] dataset archive time series classification benchmark is utilized. The datasets were selected based on their completeness, since some datasets include time series instances containing missing values. To ensure that an adequate amount of data remained for both the training and testing phases, these datasets were excluded, resulting in total of 117 datasets used.
In terms of size, the datasets vary significantly, with some containing 16 time series instances, while others encompassed up to 8,926 instances. The input lengths of these time series also showed considerable diversity, ranging from 15 to 2,844 time steps per instance in different datasets. Furthermore, the datasets included varying numbers of classes for classification tasks, with scenarios ranging from binary classifications with 2 classes to multiclass problems with up to 60 distinct classes.
The first step in data preprocessing was normalization, using ‘StandardScaler‘ from the scikit-learn library. Each time series $X$ in the dataset $D$ was standardized by applying the transformation $x_{t,\text{scaled}}=\frac{x_{t}-\mu_{t}}{\sigma_{t}}$, where $\mu_{t}$ and $\sigma_{t}$ are the mean and standard deviation for the value at time $t$ across all time series in $D$. This ensures that each time point $t$ has a mean of 0 and a standard deviation of 1, allowing for consistent scaling across the dataset. This normalization is crucial for ensuring that the input features are on a comparable scale, which helps in efficient training.
IV-B Training setup
For training, each dataset was partitioned into into a training dataset and a validation dataset in an 80:20 ratio. The test datasets are provided separately by default.
The training process across all models used the Adam optimizer with a learning rate of $0.001$. However, instead of the mean absolute error (MAE) loss function used for regression tasks in the original work, we employed the cross-entropy loss function to address the classification nature of our problem. Additionally, we introduced $L1$ regularization with a weight factor of $0.1$.
Similarly, in the MLP models, we used the same Adam optimizer, learning rate ($0.001$), and cross-entropy loss function as in the KAN models, for the same reasons. However, we applied L2 regularization with a weight factor of 1, selected through empirical tuning, instead of L1 regularization.
To ensure robustness and mitigate the effects of random initialization, each model was trained on 5 different random seeds. The results presented are averages of these 5 runs. The models were trained for 500 epochs using batch size 16.
The original KAN library [14] was modified to enable GPU support, allowing the experiments to leverage the computational power of GPUs rather than relying solely on CPU execution. Thus, all evaluations involving KAN models were conducted on an NVIDIA A100 80GB GPU.
For performance evaluation, we use precision, recall and F1 score. The precision measures how many instances predicted as a certain class actually belong to that class, expressed as: $\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$, where TP (true positives) represents the number of correctly identified positive instances, and FP (false positives) refers to the number of instances incorrectly identified as positive. Recall measures how many instances of some class were correctly detected, given by: $\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$, where FN (false negatives) represents the number of actual positive instances that the model failed to identify as positive. The F1 score is expressed as: $F1=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$, where higher values indicate better performance in balancing precision and recall (better classification).
IV-C Reference architecture configurations
Initially, we analyse the performance of KAN and MLP architectures originally proposed for forecasting by [23]. The goal is to gain insights into the generalization capabilities and transfer potential of these architectures from regression to classification problems.
As discussed in Section III, we employ two different models: KAN and MLP. Alongside the official implementation of KAN 111github.com/KindXiaoming/pykan, we also include its variant, Efficient KAN 222github.com/Blealtan/efficient-kan, which optimizes memory by reformulating the activation process, applying B-splines directly to inputs and combining them linearly, avoiding the need for large tensor expansions. It replaces input-based L1 regularization with weight-based L1, improving efficiency. The L1 regularization is now computed as mean absolute value of the spline weights [30].
KAN and Efficient KAN are configured as [$T$, 40, 40, $C$], representing a 3-depth network, where $T$ represents the length of the input time series, while $C$ denotes number of classes for classification, as per Section III. A 4-depth variant includes an additional hidden layer with 40 nodes. We retained the spline order $k=3$ and a fixed grid size $G=5$.
Similarly, the MLP architecture is configured as [$T$, 300, 300, 300, $C$], representing a 3-depth network. The 4-depth variant adds a hidden layer of $300$ nodes.
IV-D Hyperparameter impact analysis
To provide deeper insights into the KAN architecture’s performance, we conduct an analysis of hyperparameters and configuration parameters with the goal of finding the best model for time series classification on the UCR behcmark. We conducted a series of hyperparameter impact analyses by varying key hyperparameters, such as grid size $G$, depth (number of layers) $L$, and layer size (number of nodes per layer) $M$. For each of these evaluations, the models were trained using different learning rates $lr\in\{0.0001,0.001,0.01,0.1,1\}$.
The first evaluation examined the effect of grid size, where $G\in\{3,5,10,15,20\}$ while keeping depth and layer size parameters fixed. Specifically, $L=3,M=40$.
In the second evaluation, we varied the depth of the network by adjusting the number of layers $L$ between 2 and 10, with constant grid $G=5$ and layer size $M=40$.
In the final evaluation, the number of nodes per layer $M\in\{5,10,\dots 100\}$ was varied, while keeping the depth fixed at $L=3$ and grid size at $G=5$.
IV-E Complexity analysis
We evaluated resource consumption by calculating the average FLOPs over 117 datasets for one prediction, and the theoretical energy consumption (TEC) per prediction. FLOPs were computed theoretically using Eq. 8 and 11 outlined in Section III. As per Section IV-A, $T\in\{16,\ldots,8,926\}$, $C\in\{2,\ldots,60\}$.
We estimated TEC using the Eq. $\text{TEC}=\frac{\text{FLOPs}}{\text{FLOPS/Watt}}$ [31], where FLOPS/Watt represents the number of floating-point operations executed per second per watt. As the experiments were conducted on an NVIDIA A100 80GB PCIe GPU, its theoretical FLOPS/Wat of 65 GFLOPS, for float32 operations, was used for calculation.
IV-F Interpretability analysis
We demonstrate the interpretability of KANs on architecture of shape [15, 15, 15, 3], that is amenable to visualization, with a grid size G=5, learning rate lr=1, and L1 regularization of 0.01. The architecture and hyperparameters were iteratively optimized to ensure the model’s performance was competitive with MLPs and HiveCote 2.0, ensuring that the model achieved comparable results while being significantly less computationally complex. The interpretability results are showcased on the SmoothSubspace dataset from the UCR repository, which has a length of 15 and 3 output classes. However, the methods are equally effective across all datasets in the UCR repository.
Furthermore, to gain further insight into the decision making of KAN, global SHapley Additive exPlanations (SHAP) values [32] were used. SHAP offers a model-agnostic framework for global interpretability by quantifying the average contribution of each feature to the model’s predictions across the entire dataset. The contribution of a feature $x_{i}$ is represented by its Shapley value, $\phi_{i}$, which is computed based on cooperative game theory. For a model the prediction for an instance $X$ is expressed as:
$SHAP=\phi_{0}+\sum_{i=1}^{M}\phi_{i}$ (12)
where $\phi_{0}$ is the model’s baseline expected value across the dataset, and $\phi_{i}$ is the Shapley value of feature $x_{i}$, and M represents the total number of features. By aggregating these values SHAP enables insights into the global decisions of the model. In addition SHAP is also utilized to compare the feature importance of KANs to the MLP architecture.
Model Configuration Precision Recall F1 score Training Time (s)
Mean StdDev Mean StdDev Mean StdDev
MLP (3-depth) [300, 300, 300] 0.73 0.20 0.66 0.23 0.64 0.25 109.50
MLP (4-depth) [300, 300, 300, 300] 0.73 0.20 0.65 0.20 0.62 0.26 105.53
KAN (3-depth) [40, 40], $G=5$ 0.38 0.17 0.33 0.21 0.30 0.21 44.99
KAN (4-depth) [40, 40, 40], $G=5$ 0.37 0.11 0.32 0.20 0.29 0.20 56.88
Efficient KAN (3-depth) [40, 40], $G=5$ 0.72 0.20 0.70 0.21 0.70 0.22 90.80
Efficient KAN (4-depth) [40, 40, 40], $G=5$ 0.71 0.20 0.70 0.22 0.69 0.22 92.71
TABLE I: Performance of reference architectures introduced in [23] for forecasting.
V Results
This section presents a comprehensive analysis of the results through classification performance of reference architectures, computational complexity comparison, hyperparameter impact analysis, and interpretability evaluation.
V-A Classification performance analysis of reference architectures
Table I presents the results of the classifiers. The table is divided into rows representing different models and their configurations, with the first two rows listing the results for MLP models and the subsequent four rows listing the results for KAN and Efficient KAN models. The first column lists the type of model, while the second column details the specific configuration of each model, including the nodes per layer and grid size for KAN models. The following columns provide the performance metrics for each model, split into three main categories: precision, recall, and F1 score. Each of these categories is further divided into two subcolumns that show the mean and standard deviation (StdDev) of the results across multiple runs. The final column presents average time, measured in seconds, for training one model.
Looking at the first two rows of the table, the MLP models demonstrate relatively high and consistent performance across precision, recall, and F1 score metrics in both depths. For the 3-depth MLP, the mean precision is 0.77 with a standard deviation of 0.18, while the 4-depth model exhibits a similar precision mean of 0.76 and a slightly lower standard deviation of 0.16. In terms of recall, both depths show identical mean values of 0.64 and a standard deviation of 0.26, indicating no change in the model’s ability to identify relevant instances with increased depth. Additionally, both models maintain the same F1 score mean of 0.62, although the standard deviation slightly increases from 0.26 in the 3-depth model to 0.27 in the 4-depth model, suggesting a minor decrease in consistency for the F1 score with the deeper model.
Rows 3 and 4 for the original KAN implementation show a substantial drop in performance compared to the MLP models. The 3-depth KAN has a mean precision of 0.38, with a recall of 0.33 and an F1 score of 0.30. The 4-depth KAN model continues this trend with slightly lower performance, showing a precision of 0.37, recall of 0.32, and F1 score of 0.29. The further decline in performance suggests that adding depth does not improve generalization but may instead lead to overfitting, especially across multiple diverse datasets. These results indicate that KAN struggles to match the classification performance of the MLP. While previous research demonstrated that KAN performed well for time series prediction and exhibited strong generalization abilities [23], here KAN’s transferability to a different problem, classification, appears less successful.
Lower performance results come with a significantly reduced training time of $\sim 45$ seconds for 3-depth, due to the GPU acceleration, as described in Section IV-E, suggesting a trade-off between performance and efficiency. However, the training time in 4-depth increases to $\sim 57$ seconds, reflecting the impact of the added depth on computational cost.
Model Configuration Lernable Parameters Theor. FLOPs TEC (Joules)
MLP (3-depth) [300, 300, 300] 344 518 688 420 $\boldsymbol{1.008\times 10^{-5}}$
MLP (4-depth) [300, 300, 300, 300] 434 818 868 720 $1.336\times 10^{-5}$
KAN (3-depth) [40, 40], $G=5$ 257 649 6 146 993 $9.457\times 10^{-5}$
KAN (4-depth) [40, 40, 40], $G=5$ 275 289 6 566 993 $10.103\times 10^{-5}$
Efficient KAN (3-depth) [40, 40], $G=5$ 257 649 6 146 993 $9.457\times 10^{-5}$
Efficient KAN (4-depth) [40, 40, 40], $G=5$ 275 289 6 566 993 $10.103\times 10^{-5}$
TABLE II: Comparison of performance and computational characteristics across the reference models
The last two rows show results for the models trained with the Efficient KAN implementation, which offer a notable improvement in performance compared to the original KAN implementation. The Efficient KAN with 3-depth, demonstrates a significant boost in performance with a precision of 0.72, recall of 0.70, and F1 score of 0.70. Similarly, the 4-depth Efficient KAN, maintains high performance metrics with a precision of 0.71, recall of 0.70, and F1 score of 0.69. Improvements in comparison to KAN can be attributed to implementation changes, as discussed in Section IV-C.
The precision of Efficient KAN is slightly lower than that of the MLP models, which means that the model may produce more false positives. However, the Efficient KAN model compensates for this with higher recall, indicating that it is better at identifying true positives. As a result, the F1 score is also higher for Efficient KAN. This suggests that while Efficient KAN sacrifices a bit of precision, it achieves a better overall performance by improving recall and providing a more balanced model for classification. Along with the increased performance, Efficient KAN maintains fast training times. The training time for the 3-depth model is approximately 91 seconds, with a slight increase to 93 seconds for the 4-depth model. This shows that while the original KAN implementation is less successful in transferring from time series prediction to classification tasks, the Efficient KAN not only achieves this transfer effectively but also generalizes better across different datasets, making it a more effective model overall.
V-B Complexity analysis for the reference models
Table II presents the computational complexity analysis for the MLP and KAN configurations. The first two columns, listing the model and configuration, are referenced from Table I. The subsequent columns provide key complexity metrics: the third column displays the total number of lernable parameters in each model, while the fourth column reports the average theoretical FLOPs for one prediction, based on model architecture discussed in Section III, across all used UCR datasets. The last column presents the Theoretical Energy Consumption (TEC) in Joules, calculated based on the FLOPs and GPU efficiency, as per Section IV-E.
Examining the first two rows of the table, we observe that MLPs, with approximately $344k$ and $434k$ learnable parameters for depths $3$ and $4$ respectively, are the most computationally demanding models in terms of learnable parameters and training duration. Their theoretical FLOPs reach $688,420$ and $868,720$, while requiring $1.008\times 10^{-5}$ and $1.336\times 10^{-5}$ Joules of energy per single prediction, positioning them as the best-performing models in theoretical FLOPs and TEC (Total Energy Consumption).
Rows 3-4 of the table reveal that the original KAN implementation has noticeably fewer learnable parameters compared to the MLP models, consistent with the theoretical analysis in Section III-A. Despite the reduction in parameters, they come with significantly higher theoretical FLOPs, with values of $6,146,993$ for depth 3 and $6,566,993$ for depth 4, due to B-spline computation, as per Eq. 8. The energy required for a single prediction also reflects this increase, amounting to $9.457\times 10^{-5}$ Joules for depth 3 and $10.103\times 10^{-5}$ Joules for depth 4. These results illustrate a trade-off, where KAN achieves a smaller parameter count but requires greater computational resources in terms of FLOPs and energy per prediction.
The last two rows for the Efficient KAN implementation show the same number of weights and theoretical FLOPs as the original implementation, as the model architectures are identical. Despite higher energy consumption, Efficient KAN maintains fast training times. Along with the discussed increased performance, this balance of high performance and lower computational costs makes these models strong candidates for applications requiring both efficiency and effectiveness.
However, despite these improvements, MLP models remain far more energy-efficient, requiring 10 times less energy (TEC) per prediction than both KAN and Efficient KAN models.
Figure 4: Grid impact on KAN and Efficient KAN models
Figure 5: Depth impact on KAN and Efficient KAN models
Figure 6: Layer size impact on KAN and Efficient KAN models
V-C Hyperparameter impact analysis
We analyze the impact of key hyperparameters on KAN and Efficient KAN models, with each Fig. 4 5 6 illustrating results for specific hyperparameter variations as detailed in Section IV-D. Each figure includes line plots for five different learning rates, where the same color represents the results with the same learning rate for both models, with KAN depicted with solid lines and Efficient KAN with dashed lines.
The impact of grid size $G$ on model performance was evaluated across various learning rates, focusing on mean F1 score outcomes for both KAN and Efficient KAN. As it can be seen in Fig. 4, results show a general trend of decreasing F1 performance with increased grid size, which aligns with findings that larger grid configurations can lead to optimization challenges in KAN models [25]. This also suggests that while increasing grid size makes B-splines locally accurate, it potentially reduces the model’s overall performance by making it globally inaccurate. While both models’ performances decrease with grid size, Efficient KAN maintains greater stability across grid configurations and learning rates, with lower learning rates proving more effective in preserving model performance.
At the lowest learning rates, Efficient KAN achieved the highest F1 scores, with $\text{lr}=0.0001$ and $\text{lr}=0.001$ yielding nearly identical and stable results across grid sizes. For these rates, the F1 score exhibited a slight, linear decrease from $0.69$ at $G=3$ to approximately $0.66$ at $G=20$, suggesting minimal sensitivity to grid variation. The learning rate $\text{lr}=0.01$ yielded slightly lower performance than $\text{lr}=0.0001$ and $\text{lr}=0.001$, following a similarly gradual decrease across grid sizes. In contrast, higher learning rates ($\text{lr}=0.1$ and $\text{lr}=1$) resulted in significantly lower F1 scores, averaging around $0.3$ across all grid sizes, showing limited grid sensitivity but almost 50% reduced effectiveness relative to lower rates.
KAN, in contrast, exhibited an opposing learning rate trend, with performance improving as learning rates increased. All learning rates for KAN showed a similar pattern: a substantial drop in F1 score from $G=3$ to $G=10$, followed by a more gradual decline with increasing grid size. Notably, even KAN’s best configurations fell well below the best results of Efficient KAN, underscoring a clear performance gap. A deviation in KAN’s trend was seen at $\text{lr}=1$ for $G=3$, where the F1 score was unexpectedly lower than at $G=5$, unlike other rates where $G=3$ consistently achieved the highest performance.
The effect of model depth reveals similar trends, as shown in Fig. 5. Efficient KAN achieved the highest performance at the smallest learning rates ($lr=0.0001$ and $lr=0.001$), though differences between these rates became more pronounced with increasing depth. At a learning rate of $lr=0.01$, Efficient KAN’s performance declined more sharply with increasing depth, achieving a peak F1 score of $0.69$ at depth 2, but dropping to $0.64$ at depth 10. This pattern aligns with previous research indicating that higher learning rates can destabilize training by causing the model to overshoot optimal parameter values. In time-series tasks, lower learning rates in combination with Efficient KAN’s regularization mechanisms enable it to adjust gradually, allowing the model to learn nuanced temporal patterns without abrupt parameter updates, thereby supporting stable and high performance.
In contrast, KAN’s behavior across learning rates mirrored the patterns observed in the grid size analysis, with generally lower performance compared to Efficient KAN. For shallower depths (2–3), KAN achieved its best performance with higher learning rates; however, these rates led to a sharp performance drop as depth increased, with F1 scores falling below 0.2. Lower learning rates yielded a more gradual decline, stabilizing around 0.3 before slightly decreasing further. Regardless of the learning rate, increasing depth consistently resulted in lower performance. These results suggest that KAN performs best with smaller depths and benefits less from deeper configurations due to challenges in capturing complex time-series dependencies effectively at greater depths.
We further analyzed the effect of layer size, defined as the number of nodes in each layer, on model performance. Results are presented in Fig. 6. For Efficient KAN, lower learning rates ($lr=0.0001$ and $lr=0.001$) consistently yielded the best results, with $lr=0.01$ performing slightly worse. Larger learning rates led to a significant performance drop, mirroring previous observations. Among layer sizes, configurations with 5 and 10 nodes per layer generally underperformed compared to larger sizes, which yielded stable F1 scores. These smaller configurations lack the complexity required to model diverse time-series data effectively, especially given the shortest series length of 15 data points. When layer sizes exceeded this threshold, Efficient KAN maintained consistent performance by avoiding overgeneralization and better capturing the data’s intricacies.
Model Configuration Precision Recall F1 score Time (seconds)
Mean StdDev Mean StdDev Mean StdDev Mean StdDev
KAN [40], $G=5$ 0.58 0.19 0.56 0.21 0.54 0.22 6.36 2.8
Efficient KAN [40, 40], $G=3$ 0.72 0.20 0.71 0.21 0.70 0.23 95.00 205.1
MLP (3-depth) [300, 300, 300] 0.73 0.20 0.66 0.23 0.64 0.25 109.5 368.5
Hive-Cote 2.0 / 0.87 0.13 0.85 0.16 0.85 0.17 14 700.32 57 228.95
TABLE III: Comparison of top-performing KAN models vs the reference MLP and Hive-Cote2 SotA.
For KAN, the trend remained opposite: higher learning rates produced better F1 scores, with $lr=0.1$ consistently yielding the highest performance. Unlike Efficient KAN, KAN’s performance gradually improved with increasing layer size, indicating that a larger number of nodes allowed it to capture more complex dependencies in the time-series data. A notable increase in F1 performance occurred with a layer configuration of 30 nodes per layer and a learning rate of $lr=0.001$, highlighting this setup’s capacity to leverage both the model’s structure and time-series complexity. However, the worst-performing configuration for KAN was a high learning rate with very few nodes (5–20 per layer), which led to significant overgeneralization and poor performance. Smaller configurations likely struggle to represent complex patterns effectively, leading to overly simplistic models that underfit the data’s temporal dependencies.
In summary, the analyses of grid size, depth, and layer size reveal clear distinctions in how KAN and Efficient KAN respond to different hyperparameter configurations. Efficient KAN consistently performs best with smaller learning rates, showing stability across grid, depth, and layer size variations due to its inherent regularization mechanisms. This stability allows Efficient KAN to avoid overfitting and adapt effectively even with moderate layer sizes, particularly when layer sizes exceed the shortest time series length. In contrast, KAN’s optimal performance relies on higher learning rates and increased layer sizes, as these configurations enable it to capture more complex dependencies in the time-series data. KAN’s performance improves gradually with layer size and depth but is more sensitive to increases in grid size. Overall, Efficient KAN demonstrates a robust ability to maintain high performance with conservative hyperparameter settings, whereas KAN benefits more from expanded configurations to reach its optimal performance.
Out of all computed results across various hyperparameter configurations, we selected the best-performing KAN and Efficient KAN models. Their results are presented in Table III. Table is structured as Table I, along with additional column showing average training time in seconds.
Figure 7: Critical diagram of F1 score for five models
For comparison, we also include results from the MLP model discussed in Section V-A and Hive-Cote 2.0 as a state-of-the-art benchmark. Efficient KAN demonstrates comparable performance to MLP, with higher F1 scores and significantly faster training times. Although Hive-Cote 2.0 achieves the highest overall performance, Efficient KAN provides a balanced tradeoff with substantially reduced training time, making it a practical alternative for time-sensitive applications. This is further illustrated in the critical difference diagrams in Fig. 7, where Efficient KAN ranks above MLP but still below Hive-Cote 2.0, reflecting its competitive accuracy while maintaining efficiency. In contrast, KAN ranks lowest, showing a more significant deviation in performance.
V-D Interpretability analysis
[…]
Figure 8: KAN model
Figure 9: SHAP summary plots showing feature impacts in KAN model
Figure 10: SHAP summary plots showing feature impacts in MLP model
In this section, we evaluate the interpretability of the KAN model and compare it with the interpretability of MLP, following the methodology described in Section IV-F. The interpretability of the two models is demonstrated using the SmoothSubspace UCR dataset shown in Fig. 1. We compare the interpretability of the KAN model with the best MLP configuration (see Table III), which achieved an F1 score of almost 0.88 on the used dataset, which is similar to the fine-tuned KAN model utilised to demonstrate interpretability described in Section  IV-F.
Fig. 8 provides a visualization of the structure of the KAN model, where the thickness of the edges represents the importance of the connections based on the edge scores according to Section IV-F. The corresponding B-splines for these edges show which function is applied to the input features and provide information on how the input features are modified by the architecture. The x-axis of the enhanced B-spline in Fig. 8 represents the range of input features, while the y-axis indicates the corresponding outputs of the spline function applied to these inputs.
Looking at Fig. 8, it can be seen that features $x_{5}$, $x_{10}$, and $x_{12}$ emerge as the most influential for the models decision, based on the edge weights where thicker lines represent higher edge weight. Looking at Fig. 1 that input features $x_{5}$, $x_{10}$ are effective in differentiating class $y_{2}$ from classes $y_{1}$ and $y_{3}$, while feature $x_{12}$ further distinguishes between classes $y_{1}$ and $y_{3}$. Additionally, looking at class $y_{3}$ in Fig. 1, we can see that majority of the values within the dataset are whithin $[0,1]$ for for input feature $x_{12}$. Looking at the learned B-spline corresponding to this feature in Fig. 8 we can see that when $x_{12}$ is between 0 and 1, the spline increases the values up to 300, while the values outside of this range decrease to values between 0 and -200, diminishing their contribution. Similar observations can also be made for the input features $x_{5}$, $x_{10}$.
To further quantify feature contributions and also compare feature importance of KAN to MLP, we apply SHAP analysis as described in Section IV-F. Fig. 9 and 10 display the feature importance for each class of the SmoothSubspace dataset for both KAN and MLP, sorted from the most important to the least important feature for each predicted class. Each figure is divided into three sections, corresponding to the three output classes. In the plots, each feature’s contribution is represented by its SHAP value, which indicates whether it positively contributes to the prediction or negatively. Feature values are color-coded, with red indicating high values and blue low values. As it can be seen in Fig. 1, $2$ represents the highest value, while $-2$ represents the lowest value. Middle-range values between $-1$ to $1$, are shown as purple dots. Purple dots correspond to the flat parts of the series, where consecutive values show minimal variation, while red and blue dots are associated with the highly fluctuating sections of the series.
Examining the SHAP plot for the KAN model in Fig. 9, it can be seen that the top contributors, features $x_{10}$ and $x_{12}$ consistently stand out, aligning with KAN’s interpretability findings from Fig. 8. Interestingly, compared to the feature importance from Fig. 8, feature $x_{5}$ is not present among the top contributors according to SHAP. This contrast shows how different interpretability methods can yield complementary insights, each capturing unique aspects of feature importance.
Looking at Fig. 10, the SHAP feature importance scores are presented for the MLP. As it can be seen, feature $x_{7}$ is the most important in distinguishing classes $y_{1}$ and $y_{2}$, compared to $x_{10}$ for KAN model. However, for both MLP and KAN the $x_{10}$ and $x_{7}$ are shown as top 4 contributing features in distinguishing between the classes, indicating that both MLP and KAN model in general consider similar features as key to distinguish between the three classes.
VI Conclusion
The aim of this study was to conduct a comprehensive and robust exploration of the KAN architecture for time series classification. To achieve that, we first investigated the transferability of reference Kolmogorov-Arnold Networks (KANs) from regression to classification tasks followed by key hyperparameter search on two KAN implementations in view of finding the best architecture for the task at hand and concluding with an interpretability analysis. Our findings show that the EfficientKAN implementation has a significant performance improvement over the original KAN and is also superior to MLPs by 6 percentage points in F1 score on the reference architectures. The hyperparameter analysis revealed that Efficient KAN consistently outperformed KAN across variations in grid size, depth, and layer size, showcasing its robustness and adaptability to diverse datasets. Efficient KAN proved more stable than KAN across grid sizes, depths, and layer configurations, particularly with lower learning rates. The interpretability of KANs was confirmed through SHAP analysis and B-spline visualization, highlighting their potential for transparent decision-making. In contrast, MLPs, though competitive in performance, exhibited less interpretability. Future work could explore hybrid architectures and further optimization to enhance KAN’s application scope.
Acknowledgments
This work was supported by the Slovenian Research Agency (Javna Agencija za Raziskovalno Dejavnost RS) under grant P2-0016. A preprint has previously been published [33].